# Debiased Machine Learning and Network Cohesion for Doubly-Robust Differential Reward Models in Contextual Bandits

### v2

@@ 3.1 DML and Doubly Robust Differential Reward
Both $\mathbb{E}_{p(s,\bar{a})}\left[(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a}))^{2}\right]=o_{P}(k^{-1/2})$ and $\mathbb{E}_{p(s)}\left[(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}\right]=o_{P}(k^{-1/2})$. Further, $|\hat{f}_{i,t}|\leq 2B$.
@@
There exists $c>0$ such that $\tilde{\sigma}_{i,t}^{2}>c$ for all $i,t$.
3.2 Nearest Neighbor Regularization
Above, we considered a single individual $i$ under a time-invariant linear differential reward function; i.e., $\boldsymbol{x}(s_{i,t},a_{i,t})^{\top}\theta$ where $\theta\in\mathbb{R}^{p}$. Here, we consider the setting of $N$ independent individuals and a time-invariant linear differential reward with an individual-specific parameter; i.e., $\theta_{i}\in\mathbb{R}^{p}$. If $\theta_{i}$ were known a priori, then one could construct a network based on $L_{2}$-distances $\{d(i,j):=\|\theta_{i}-\theta_{j}\|_{2}^{2}\}_{j\not=i}$.
Specifically, define a graph $G=(V,E)$ where each user represents a node, e.g., $V:=[N]$, and $(i,j)$ is in the edge set $E$ for the smallest $M\ll N$ distances. The working assumption is that connected users share similar underlying vectors $\theta_{i}$, implying that the rewards received from one user can provide valuable insights into the behavior of other connected users. Mathematically, $(i,j)\in E$ implies that $\|\theta_{i}-\theta_{j}\|$ is small.
We define the Laplacian via the $N\times MN$ incidence matrix $B$. The element $B_{v,e}$ corresponds to the $v$-th vertex (user) and $e$-th edge. Denote the vertices of $e$ as $v_{i}$ and $v_{j}$ with $i>j$. $B_{v,e}$ is then equal to 1 if $v=v_{i}$, -1 if $v=v_{j}$, and 0 otherwise. The Laplacian matrix is then defined as $L=BB^{\top}$. We can then adapt (4) by summing over participants and including a network cohesion penalty similar to Yang et al. (2020):
$\text{tr}(\Theta^{\top}L\Theta)=\sum_{(i,j)\in E}\|\theta_{i}-\theta_{j}\|_{2}^{2},$
where $\Theta:=(\theta_{1},\ldots,\theta_{N})^{\top}\in\mathbb{R}^{N\times p}$. The penalty is small when $\theta_{i}$ and $\theta_{j}$ are close for connected users. Following Assumption 2 and the above discussion, we further assume:
@@ 3.2 Nearest Neighbor Regularization
There exists $D\in\mathbb{R}^{+}$ such that $\|\theta_{i}-\theta_{j}\|_{2}^{2}\leq D$, $\forall i,j$, and $D$ is known.
4 DML Thompson Sampling with Nearest Neighbor Regularization
4.1 Algorithm
Based on Section 3, we can now formally state our proposed DML Thompson Sampling with Nearest Neighbor Regularization (DML-TS-NNR) algorithm. In our study, we adopt a sequential recruitment setting in which individuals’ enrollment occurs in a staggered manner to mimic the recruitment process in real mHealth studies. More specifically, we first observe individual $i=1$ at time $t=1$. Then we observe individuals $i=(1,2)$ at times $t=(2,1)$. After $k$ time steps, we observe individuals $i\in[k]$ at times $(k+1-i,k-i,\ldots,1)$ respectively. We then observe these individuals in a random sequence one at a time before moving to stage $k+1$. Define $\mathcal{O}_{k}=\{(i,t):i\leq k\>\&\>t\leq k+1-i\}$ be the set of observed time points across all individuals at stage $k$. Again see Figure 3 in Appendix A for a visualization.
By performing a joint asymptotic analysis with respect to the total number of individuals ($N$) and time points ($T$), we can relax the assumption of a single time-invariant linear advantage function and allow $\theta_{i,t}\in\mathbb{R}^{p}$ to depend on both the individual $i$ and time $t$. Here, we let $\theta_{i,t}=\theta+\theta_{i}^{\textrm{ind}}+\theta_{t}^{\textrm{time}}$; i.e., we include (i) an individual-specific, time-invariant term $\theta_{i}$, and (ii) a shared, time-specific term $\theta_{t}$. This setup is similar to the intelligentpooling method of Tomkins et al. (2021); however, rather than assume individuals and time points are unrelated i.i.d. samples, we assume knowledge of some network information (e.g., the similarity of certain individuals or proximity in time) and regularize these parameters accordingly to ensure network cohesion.
The DML-TS-NNR algorithm is shown in Algorithm 1. To see the motivation, consider the following. We first assume that we have access to two nearest neighbor graphs, $G_{\text{user}}$ and $G_{\text{time}}$, where each characterizes proximity in the user- and time-domains, respectively. Then at stage $k$, we estimate all parameters, e.g. $\Theta_{k}=\textrm{vec}[\left(\theta,\theta_{1}^{\textrm{ind}},\ldots,\theta_{k}^{\textrm{ind}},\theta_{1}^{\textrm{time}},\ldots,\theta_{k}^{\textrm{time}}\right)]\in\mathbb{R}^{p(2k+1)}$ , by minimizing the following penalized loss function $L_{k}(\Theta_{k};\lambda,\gamma)$, which is defined as the following expression:
$\begin{split}&\sum_{(i,t)\in\mathcal{O}_{k}}\tilde{\sigma}_{i,t}^{2}\left(\tilde{R}_{i,t}^{\hat{f}^{(m)}}-\boldsymbol{x}(S_{i,t},A_{i,t})^{\top}(\theta+\theta_{i}^{\textrm{ind}}+\theta_{t}^{\textrm{time}})\right)^{2}+\\ &\gamma\left(\|\theta\|_{2}^{2}+\sum_{i=1}^{k}\|\theta_{i}^{\textrm{ind}}\|_{2}^{2}+\sum_{t=1}^{k}\|\theta_{t}^{\textrm{time}}\|_{2}^{2}\right)+\lambda\left(\text{tr}\left(\Theta_{\text{user}}^{\top}L_{\text{user}}\Theta_{\text{user}}\right)+\text{tr}\left(\Theta_{\text{time}}^{\top}L_{\text{time}}\Theta_{\text{time}}\right)\right),\end{split}$ (5)
where $\Theta_{\textrm{user}},\Theta_{\textrm{time}}\in\mathbb{R}^{p\times k}$. In comparison to existing methods, the primary novelty in Equation (5) is that (1) the observed outcome $R_{i,t}$ is replaced by a pseudo-outcome $\tilde{R}^{f^{(k)}}_{i,t}$ and (2) the doubly-robust pseudo-outcome leads to a weighted least-squares loss with weights $\tilde{\sigma}_{i,t}^{2}$. The network cohesion penalties and time-specific parameters have been considered elsewhere (Yang et al., 2020; Tomkins et al., 2021), though, not together. For more details regarding Algorithm 1, please refer to Appendix B.
Algorithm 1 DML-TS with Nearest Neighbor Regularization (DML-TS-NNR)
  Input: $\delta$, $\sigma$, $c$, $C$, $m$, $\lambda$, $\gamma$, $L$, $B_{w}$, $D_{w}$
  Set $L_{\otimes}=L\otimes I_{p}$ and $B=k\frac{\lambda}{\sqrt{\gamma}}\left(D_{\text{ind}}+D_{\text{time}}\right)+\sqrt{\gamma n}\left(B_{\text{ind}}+B_{\text{time}}\right)$
@@ 4.1 Algorithm
  Initialize: $V_{0}=\operatorname{diag}(\gamma I_{p},\lambda L^{\text{ind}}_{\otimes}+\gamma I_{kp},\lambda L^{\text{time}}_{\otimes}+\gamma I_{kp})$ and $b_{0}=\mathbf{0}$
@@
     Option 1: Randomly assign $(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}$ to one of the $M$ partitions
     Observe Context variable $S_{l}=S_{i_{l},k+i_{l}-1}$
     Set $\hat{\Theta}_{k}=V_{k}^{-1}b_{k}$
     Calculate
where $v_{k}^{2}$ is defined in Corollary 4
     Generate $\eta_{k}\sim\mathcal{D}^{TS}$ and compute
$\tilde{\Theta}_{k}=\hat{\Theta}_{k}+\beta_{k}(\delta^{\prime})V_{k}^{-1/2}\eta_{k}$
     For each $(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}$ select $A_{i,t}$ that maximizes:
$\boldsymbol{x}(S_{i,t},a)^{\top}\left(\tilde{\theta}+\tilde{\theta}_{i}^{\text{ind}}+\tilde{\theta}_{t}^{\text{time}}\right)$
     Observe rewards $R_{i,t}$
     Construct feature $\boldsymbol{x}_{t}=\boldsymbol{x}(S_{t},A_{t})$ and $\phi_{i,t}=\phi(\boldsymbol{x}_{i,t})$
     Option 1: Re-construct predictions for all $\hat{f}^{(m)}$ partitions for $m=1,\ldots,M$ and re-compute all pseudo-outcomes $\tilde{R}_{i,t}^{\hat{f}^{(m)}}$ for all $(i,t)\in\mathcal{O}_{k}$.
     Option 2: Construct predictions for next stage $\hat{f}^{(k)}$ partitions and compute pseudo-outcomes $\tilde{R}_{i,t}^{\hat{f}^{(k)}}$ only for those $(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}$
     Update $V_{k}=V_{k-1}+\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\tilde{\sigma}^{2}_{i,t}\phi_{i,t}\phi_{i,t}^{\top}$ and $b_{k}=b_{k-1}+\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\tilde{\sigma}_{i,t}^{2}\tilde{R}_{i,t}^{\hat{f}}\phi_{i,t}$
@@
4.2 Regret Analysis
Given the knowledge of true parameters $\Theta$, the optimal policy is simply to select, at decision time $t$ for individual $i$, the action $a^{*}_{i,t}=\arg\max_{a\in\mathcal{A}}\boldsymbol{x}(S_{i,t},a)^{\top}(\theta+\theta_{i}+\theta_{t})$ given the state variable $S_{i,t}$. This leads us to evaluate the algorithm by comparing it to this optimal policy after each stage. Given both the number of individuals and the number of time points increases per stage, we define stage $k$ regret to be the average across all individuals at stage $k$:
$\boldsymbol{\operatorname{Regret}}_{K}=\sum_{k=1}^{K}\frac{1}{k}\left\{\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\left[\boldsymbol{x}(S_{i,t},a_{i,t}^{*})^{\top}(\theta^{*}+\theta_{i}^{\star}+\theta_{t}^{\star})-\boldsymbol{x}(S_{i,t},A_{i,t})^{\top}(\theta^{*}+\theta_{i}^{\star}+\theta_{t}^{\star})\right]\right\}$
This regret is a version of pseudo-regret Audibert et al. (2003). Compared to standard regret, the randomness of the pseudo-regret is due to $\{A_{i,t}\}_{t=1}^{\top}$ since the error terms $\{\epsilon_{i,t}\}_{t=1}^{\top}$ are removed in the definition.

### v3

@@ 4.1 Debiased Machine Learning
The working model, $f_{i,t}$, satisfies $\mathbb{E}_{p(s,\bar{a})}\left[(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a}))^{2}\right]=o_{P}(k^{-1/2})$ and $\mathbb{E}_{p(s)}\left[(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}\right]=o_{P}(k^{-1/2})$.
Examples of supervised learning methods that converge at this rate (or faster) include parametric models of fixed dimension, local mean smoothers with small dimension (Wasserman, 2006, Chapter 5), and random forests and feedforward neural networks under certain regularity conditions (Biau, 2012; Farrell et al., 2021). Because of the Neyman Orthogonality, we can achieve sub-linear regret even if $f_{i,t}$ is misspecified. Assumption 3 guarantees better asymptotic variance than action-centering (AC) bandits Greenewald et al. (2017). Without it, the worst case performance (i.e., extreme model misspecification) can lead infinite asymptotic variance and perform empirically much worse than AC bandits (Bang & Robins, 2005). In mHealth contexts, we believe the assumption is fairly weak and plausible. The final requirement of independence, (c), can be accomplished via sample-splitting techniques. Below, we present two such options.
Option 1: The first option exploits the fact that outcomes across different users are independent. Consequently, we can perform sample splitting as follows. Step 1: Randomly assign each user $i$ to one of $J$ folds. Let $I_{j}$ denote the set containing user indexes for the $j$-th fold and let $I_{j}^{\complement}$ denote its complement. Step 2: For each fold, use a supervised learning algorithm to estimate the working model for $r_{i,t}(s,a)$ denoted $\hat{f}_{i,t}^{(j)}(s,a)$ using $I_{j}^{\complement}$. Step 3: Construct the pseudo-outcomes using (3).
Option 2: Because we do not expect an adversarial environment in mHealth, we may be willing to assume that the errors, $\epsilon_{i,t}$, are independent across time. In this case, we can adapt the procedure given above by randomly assigning data to folds at the level of $(i,t)$. In this case, $I_{j}$, contains the $(i,t)$ pairs assigned to fold $j$. This option is more powerful than option 1 in that it enables us to learn heterogeneity across users in the nuisance function, $f_{i,t}$, leading to greater variance reduction.
Lastly, we note two strategies that can reduce the computational demands of sample splitting. The first is to employ certain stable estimators that do not require sample splitting, e.g., bagged estimators using subsampling, as discussed in Chen et al. (2022). The second is to fit $f_{i,t}$ in an online fashion.
4.2 Estimation via Weighted Least Squares
Having formed $\tilde{R}_{i,t}^{f}(s,\bar{a})$, we now explain how to estimate the parameters, $\theta_{i,t}\in\mathbb{R}^{p}$, determining the differential reward, $\Delta_{i,t}(s,a)$. We assume that $\theta_{i,t}$ consists of three components:
@@ 4.2 Estimation via Weighted Least Squares
$\theta_{i,t}=\theta^{\text{shared}}+\theta^{\text{user}}_{i}+\theta^{\text{time}}_{t}$ for all $i,t$.
The parameter $\theta^{\text{shared}}$ is a “structural parameter” shared across all $i,t$. In contrast, $\theta^{\text{user}}_{i}$ and $\theta^{\text{time}}_{t}$ are “incidental parameters” specific to a given user and time point, respectively. We place these parameters in a single column vector, $\theta$, as $\textrm{vec}(\theta^{\text{shared}},\theta^{\text{user}}_{1},\ldots,\theta^{\text{user}}_{K},\theta^{\text{time}}_{1},\ldots,\theta^{\text{time}}_{K})$, where $K$ is the total number of stages. We then form a corresponding feature vector, $\phi_{i,t}=\phi(x_{i,t})$, where $\phi$ inserts $x_{i,t}$ into the positions corresponding to $\theta^{\text{shared}}$, $\theta^{\text{user}}_{i}$, and $\theta^{\text{time}}_{t}$ with zeros in all other locations; i.e., letting $p:=\text{dim}(x_{i,t})$, and the feature vector $\phi(x_{i,t})\in\mathbb{R}^{(2K+1)p}$ is defined as follows:
[…]
We could then estimate $\theta_{i,t}$ via a weighted least squares (WLS) regression that minimizes the following objective function:
$\ell_{\text{WLS}}(\theta)=\sum_{i=1}^{n}\sum_{t=1}^{K-i+1}\tilde{\sigma}_{i,t}^{2}\big{(}\tilde{R}_{i,t}^{f}-\phi_{i,t}^{\top}\theta\big{)}^{2},$ (5)
where $\tilde{\sigma}_{i,t}^{2}=\pi_{i,t}(0|S_{i,t})\cdot(1-\pi_{i,t}(0|S_{i,t}))$. However, $\ell_{\text{WLS}}(\theta)$ is over-parameterized and does not leverage the available network information. The next section modifies $\ell_{\text{WLS}}(\theta)$ by adding regularization, producing the final loss function used in DML-TS-NNR.
4.3 Nearest-Neighbor Regularization
We assume access to network information relating neighboring users and time points. These networks are defined by graphs $G_{\text{user}}=(V_{\text{user}},E_{\text{user}})$ and $G_{\text{time}}=(V_{\text{time}},E_{\text{time}})$, where $V$ denotes vertices and $E$, edges. In practice, these networks can be formed via known clusters (e.g., shared schools, companies, geographic locations) or similar covariates. In the case of $G_{\text{time}}$, a reasonable choice is to construct a graph of neighboring sequential time points; i.e., $t=1\leftrightarrow t=2\leftrightarrow t=3\dots$. We assume that neighboring units have similar incidental parameters as follows:
@@ 4.3 Nearest-Neighbor Regularization
There exists $D_{\text{user}},D_{\text{time}}\in\mathbb{R}^{+}$ such that $\|\theta^{\text{user}}_{i}-\theta^{\text{user}}_{j}\|_{2}^{2}\leq D_{\text{user}}$ and $\|\theta^{\text{time}}_{i}-\theta^{\text{time}}_{j}\|_{2}^{2}\leq D_{\text{time}}$, for all $i,j$.
Assumption 5 suggests we may be able to improve our estimates of $\theta$ by regularizing neighboring values of $\theta^{\text{user}}_{i}$ and $\theta^{\text{time}}_{i}$ toward each other. This can be accomplished by using an $L^{2}$ Laplacian penalty. We illustrate the idea using $G_{\text{user}}$; the application to $G_{\text{time}}$ is similar.
First, we form the incidence matrix, $Q\in\mathbb{R}^{N\times N}$, with the entry $Q_{v,e}$ corresponding to the $v$-th vertex (user) and $e$-th edge. Denote the vertices of $e$ as $v_{i}$ and $v_{j}$ with $i>j$. $Q_{v,e}$ is then equal to 1 if $v=v_{i}$, -1 if $v=v_{j}$, and 0 otherwise. The Laplacian matrix is then defined as $L=QQ^{\top}$. Similar to Yang et al. (2020), we can then form a network cohesion penalty across users as follows:
$\text{tr}(\Theta_{{\text{user}}}^{\top}L_{\text{user}}\Theta_{{\text{user}}})=\sum_{(i,j)\in E_{\text{user}}}\|\theta^{{\text{user}}}_{i}-\theta^{{\text{user}}}_{j}\|_{2}^{2},$
where $\Theta_{\text{user}}:=(\theta^{\text{user}}_{1},\ldots,\theta^{\text{user}}_{N})^{\top}\in\mathbb{R}^{N\times p}$ and $L_{\text{user}}$ is the Laplacian matrix for users. The penalty is small when $\theta^{\text{user}}_{i}$ and $\theta^{\text{user}}_{j}$ are close for connected users. We employ a shared regularization hyperparameter $\lambda$, for $G_{\text{user}}$ and $G_{\text{time}}$, and include a standard $L^{2}$ penalty on $\theta$ with hyperparameter $\gamma$. The full penalization matrix $V_{0}\in\mathbb{R}^{(2K+1)p\times(2K+1)p}$ is
@@
$V_{0}=\operatorname{diag}(\gamma I_{p},\lambda L^{\text{user}}_{\otimes}+\gamma I_{Kp},\lambda L^{\text{time}}_{\otimes}+\gamma I_{Kp}),$ (6)
where $L^{\text{user}}_{\otimes}:=L_{\text{user}}\otimes I_{p}$ , the Kronecker product of $L_{\text{user}}$ with $I_{p}$. We then adapt $\ell_{\text{WLS}}(\theta)$ to include a corresponding penalty term:
$\ell(\theta)=\sum_{i=1}^{n}\sum_{t=1}^{K-i+1}\tilde{\sigma}_{i,t}^{2}\big{(}\tilde{R}_{i,t}^{f}-\phi_{i,t}^{\top}\theta\big{)}^{2}+\theta^{\top}V_{0}\theta.$ (7)
In Section 5 we show that (7) leads to a Thompson sampling algorithm for action selection.
5 Thompson Sampling Algorithm
We now describe the Thompson sampling procedure and provide a high-probability regret bound.
5.1 Algorithm Description
The minimizer of the weighted regularized least squares loss in (7) can be expressed as $\hat{\theta}=V^{-1}b$, where
$V=V_{0}+\sum_{i=1}^{n}\sum_{t=1}^{K-i+1}\tilde{\sigma}_{i,t}^{2}\phi_{i,t}\phi_{i,t}^{\top},\quad b=\sum_{i=1}^{n}\sum_{t=1}^{K-i+1}\tilde{R}_{i,t}^{f}\sigma_{i,t}^{2}\phi_{i,t}.$
In analogy to Bayesian linear regression, the penalized and weighted Gram matrix, $V$, plays the role of a posterior precision matrix. This motivates an algorithm in which we randomly perturb the point estimate, $\hat{\theta}$, scaling by an inverse square-root matrix, $V^{-1/2}$, times a mean-zero random deviate, $\eta$. The perturbed estimate, $\tilde{\theta}$, can be written as $\tilde{\theta}=\hat{\theta}+\beta(\delta)V^{-1/2}\eta,$ where $\beta(\delta)$ is defined as
$\beta(\delta)$ $=v\left[2\log\left(\frac{\det(V)^{1/2}}{\det(V)^{1/2}\delta/2}\right)\right]^{1/2}+B,$ (8)
$B$ $=K\frac{\lambda}{\sqrt{\gamma}}\left(D_{\text{user}}+D_{\text{time}}\right)+\sqrt{\gamma n}\left(B_{\text{user}}+B_{\text{time}}\right).$ (9)
The value $v$ is defined in Corollary 4 and can be set to a sufficiently large constant. The hyperparameter $\delta$ determines the probability with which the regret bound holds. The random deviate, $\eta$, is drawn from a distribution, $\mathcal{D}^{TS}$, satisfying the technical conditions given in Definition 1 in Appendix E. In practice, we recommend setting $\mathcal{D}^{TS}$ to a multivariate Gaussian distribution or t-distribution to simplify calculations. Given $\tilde{\theta}$, we then compute $\bar{A}_{i,t}=\operatorname{argmax}_{a\in[q]\backslash\{0\}}\phi_{i,t}^{\top}\tilde{\theta}$ and randomly select $A_{i,t}=0$ with probability
$\pi_{i,t}(0|\bar{A}_{i,t}):=\max\big{(}\pi_{\min},\min\{\pi_{\max},\Pr(\phi_{i,t}^{\top}\tilde{\theta}>0)\}\big{)},$ (10)
or $A_{i,t}=\bar{A}_{i,t}$ with probability $1-\pi_{i,t}(0|\bar{A}_{i,t})$. The action selection is summarized in Algorithm 1.
Algorithm 1 Action selection
0:  $V$, $b$, $\pi_{\min}$, $\pi_{\max}$, $\phi_{i,t}$, $\beta(\delta)$
  $\hat{\theta}=V^{-1}b$
  Sample $\eta\sim\mathcal{D}^{TS}$
  $\tilde{\theta}=\hat{\theta}+\beta(\delta)V^{-1/2}\eta$, with $\beta(\delta)$ given in (8)
  $\bar{A}_{i,t}=\operatorname{argmax}_{a\in[q]\backslash\{0\}}\phi_{i,t}^{\top}\tilde{\theta}$
  Calculate $\pi_{i,t}(0|\bar{A}_{i,t})$ according to (10)
  With probability $\pi_{i,t}(0|\bar{A}_{i,t})$, play action 00; otherwise, play action $\bar{A}_{i,t}$
Having detailed the action selection algorithm, we now summarize DML-TS-NNR in Algorithm 2. The algorithm includes an outer loop that iterates through the stages from $k=1$ to $k=K$. Within each stage, we handle users sequentially, iteratively selecting actions according to Algorithm 1, observing rewards, and updating $V$ and $b$ accordingly.
Algorithm 2 DML-TS-NNR
  Initialize $V_{0}$ according to (6)
  $V\leftarrow V_{0}$
  $b\leftarrow 0$
@@ 5.1 Algorithm Description
     for $(i,j)=(1,k),(2,k-1),\dots,(k,1)$  do
        Choose an arm according to Algorithm 1 using $V$ and $b$
        Observe rewards, $R_{i,t}$
        Construct pseudo-rewards, $\tilde{R}_{i,t}^{f^{(m)}}$, as explained in Section 4.1
        Set $\phi=\phi(x_{i,t})$ as described in (4)
        Update weighted Gram matrix: $V\leftarrow V+\tilde{\sigma}^{2}_{i,t}\phi_{i,t}\phi_{i,t}^{\top}$
        Update feature-outcome products: $b\leftarrow b+\tilde{\sigma}_{i,t}^{2}\tilde{R}_{i,t}^{f}\phi_{i,t}$
@@
  end for
5.2 Regret bound
We first define regret, $\boldsymbol{\operatorname{Regret}}_{K}$, in terms of the following average across stages:
$\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}\pi_{i,t}^{\star}(\bar{A}_{i,t}^{*}|S_{i,t})\cdot x(S_{i,t},\bar{A}_{i,t}^{*})^{\top}\theta^{*}_{i,t}\\ -\pi_{i,t}(\bar{A}_{i,t}|S_{i,t})\cdot x(S_{i,t},\bar{A}_{i,t})^{\top}\theta^{*}_{i,t}\big{]}.$
The inner sum calculates regret (the difference between the expected reward of the optimal action and the chosen action) weighted by the probability of the action, following Greenewald et al. (2017), for all individuals at stage $k$. We then average these values to account for the growing number of pairs $(i,t)$ as the stages progress. The cumulative regret is the sum of these averaged values over all stages.
The rewards associated with the baseline action are excluded since $\Delta_{i,t}(S_{i,t},0)=0$. $\boldsymbol{\operatorname{Regret}}_{K}$ counts as a pseudo-regret because the randomness attributed to $\{\epsilon_{i,t}\}_{t=1}^{K}$ has been eliminated (Audibert et al., 2003). Before presenting the regret bound, we introduce one more assumption:
Assumption 6.
$\|x(s,a)\|\leq 1$ for all $s$ and $a$. Also, there exists a known value $M\in\mathbb{R}^{+}$ bounding both $\|\theta_{i,t}\|$ and $|g_{t}(s)|$ for all $i$, $t$, and $s$. Further, $|f_{i,t}|\leq 2M$.